OpenStudy (anonymous):

(a) With A = [{1, i, 0}, {i, 0, 1}], use elimination to solve Ax = 0 (b) Show that the nullspace you just computed is orthogonal to C(A^H) and not to the usual row space C(A^T). The four fundamental spaces in the complex case are N(A) and C(A) as before, and then N(AH) and C(A^H) The nullspace I got from part a) does not fit with part b). ie, it is orthogonal to C(A^T), not C(A^H). the null space I got is {i,-1,1}. Any help will be appreciated.

OpenStudy (anonymous):

Since A has 2 columns, null space solution of x can be in terms of {x1, x2}. But you got {i,-1,1} which is in terms of {x1, x2, x3]. How is that possible??

OpenStudy (helder_edwin):

I assume the matrix is \[ \large \begin{pmatrix} 1 & i\\ i & 0\\ 0 & 1 \end{pmatrix} \] which can be reduced to \[ \large \begin{pmatrix} 1 & 0\\ 0 & 1\\ 0 & 0 \end{pmatrix} \] so the null space is \[ \large \left\langle\begin{pmatrix} 0\\ 0\\ 1 \end{pmatrix}\right\rangle \]